OpenStudy (anonymous):

If 40 is increased to 56, what is the percent of change?

OpenStudy (anonymous):

first find the actual increase, which is 16

OpenStudy (anonymous):

then compute \(16\div 40\) and move the decimal two places to the right to turn the decimal in to a percent

OpenStudy (anonymous):

Percent increase =[ (56 - 40) / 40] x 100 = 16 / 40 = (2 / 5) * 100 = ? So what was done here was simply to find the actual increase from the first value of 40 to 56. This was an increase of 16. Now since we are finding the percent of change, we divide this increase by our initial value. This gives us a decimal. To turn this decimal in to a percent, we multiply it by 100. So now, can you compute the final percentage? @kidd504